# Pulsating with discrete symmetry

### v2

@@ Abstract
The discrete symmetry of PLM seems to preclude defect turbulence, at variance with the standard CGLE [30] and its recent generalization [27]. To capture such a turbulence, one could introduce energy levels which maintain the discrete nature of PLM while breaking its symmetry. The arrested phase would no longer be degenerate, thus opening the door to local excitations nucleating defects from a homogeneous phase, as reported in [27]. In this context, it would be interesting to explore how density fluctuations affect defect nucleation, at both the hydrodynamic and the particle-based levels.
@@ Acknowledgements.
We acknowledge fruitful discussions with L. K. Davis, M. Esposito, J. Meihbom, and Y. Zhang. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101056825 (NewGenActive), and from the Luxembourg National Research Fund (FNR), grant reference 14389168.

### v3

@@ Abstract
The discrete symmetry of our DOM seems to preclude a defect turbulence, at variance with the standard CGLE [35] and its recent generalization [27]. To capture such a turbulence, one could introduce energy levels which maintain the discrete nature of our DOM while breaking its symmetry. The arrested phase would no longer be degenerate, thus opening the door to local excitations nucleating defects from a homogeneous phase, as reported in [27, 33]. In this context, it would be interesting to explore how density fluctuations affect the defect nucleation at the hydrodynamic and microscopic levels [58]. Finally, our DOM can be straightforwardly adapted to account for thermodynamic consistency [59], allowing one to study the energetics of the corresponding patterns [60, 61, 62].
@@ Acknowledgements.
We acknowledge fruitful discussions with L. K. Davis, M. Esposito, J. Meihbom, and Y. Zhang. This project has received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101056825 (NewGenActive), and from the Luxembourg National Research Fund (FNR), grant reference 14389168. A.M. acknowledges financial support from the project MOCA funded by MUR PRIN2022 grant No. 2022HNW5YL